OpenStudy (anonymous):

What is the slope of the line that contains the points (-1, -1) and (3, 15)? Enter your answer as an integer.

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \rm {Slope}= \frac{(y_1-y_2)}{(x_1-x_2)} }\) in your case, \(\LARGE\color{blue}{ \frac{(-1)-(15)}{(-1)-(3)} }\)

OpenStudy (anonymous):

-4

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \frac{(-1)-(15)}{(-1)-(3)} =\frac{-16}{-4}=\underline{~~4 ~~} }\)

OpenStudy (solomonzelman):

Makes sense, right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

negative turns positive

OpenStudy (solomonzelman):

yes